# Estimation of Concept Explanations Should be Uncertainty Aware

### 13 Dec 2023 v1

@@ Abstract
Model explanations are very valuable for interpreting and debugging prediction models. We study a specific kind of global explanations called Concept Explanations, where the goal is to interpret a model using human-understandable concepts. Recent advances in multi-modal learning rekindled interest in concept explanations and led to several label-efficient proposals for estimation. However, existing estimation methods are unstable to the choice of concepts or dataset that is used for computing explanations. We observe that instability in explanations is due to high variance in point estimation of importance scores. We propose an uncertainty aware Bayesian estimation method, which readily improved reliability of the concept explanations. We demonstrate with theoretical analysis and empirical evaluation that explanations computed by our method are more reliable while also being label-efficient and faithful.
@@ 1 Introduction
With an ever increasing complexity of ML models, there is an increasing need to explain them. Concept-based explanations are a form of interpretable methods that explain predictions using high-level and semantically meaningful concepts (Kim et al., 2018). They are aligned with how humans communicate their decisions (Yeh et al., 2022) and are shown (Kim et al., 2018, 2023b) to be more preferable over explanations using salient input features (Ribeiro et al., 2016; Selvaraju et al., 2017) or salient training examples (Koh & Liang, 2017). Concept explanations also show potential in scientific discovery (Yeh et al., 2022) and for encoding task-specific prior knowledge (Yuksekgonul et al., 2022).
@@
Concept explanations explain a pretrained prediction model by estimating the importance of concepts using two human-provided resources (1) a list of potentially relevant concepts for the task, (2) a dataset of examples usually referred to as the probe-dataset. Estimation usually proceeds in two steps (a) compute the log-likelihood of concept given an example called concept activations, and (b) aggregate their local activation scores into a globally relevant explanation. For example, the concept wing is considered important if the information about the concept is encoded in all examples of the plane class in the dataset. Owing to example-agnostic and classifier-level nature of concept explanations they are easy to interpret and have witnessed wide recognition in diverse applications (Yeh et al., 2022).
@@
Despite their easy interpretation, concept explanations are known to be unreliable and data expensive. Ramaswamy et al. (2022a) showed that existing estimation methods are sensitive to the choice of concept set and dataset raising concerns over their interpretability. Another major limitation of concept-based explanation is the need for datasets with concept annotations in order to specify the concepts. Increasingly popular multi-modal models such as CLIP (Radford et al., 2021) present an exciting alternate direction to specify relevant concepts, especially for common image applications through their text description. Recent work has explored using multi-modal models for training concept-bottleneck models (Oikarinen et al., 2023; Yuksekgonul et al., 2022; Moayeri et al., 2023), but they are not yet evaluated for generating post-hoc concept explanations.
@@
Our objective is to generate reliable concept explanations without requiring datasets with concept annotations. We begin by observing that existing estimation methods do not model noise in the estimation pipeline leading to high variance and unreliable explanations. We identify at least two causes of uncertainty (Section 4.1 presents more concrete scenarios) leading to unreliable explanations (1) When a concept is missing from the probe-dataset, we cannot estimate its importance with confidence. Reporting uncertainty over estimated importance of a concept can thus help the user make a more informed interpretation, (2) When a concept is hard or irrelevant to the task their corresponding activations predicted from the representation layer of the model-to-be-explained are expected to be noisy. For example, it is harder to recognise the concept whiskers when compared with the concept wings. The noise or uncertainty in concept activations either due to their absence, hardness, or relevance if not modelled cascades into noise in explanations. Appreciating the need to model uncertainty, we present an estimator called Uncertainty-Aware Concept Explanations (U-ACE), which we show is instrumental in improving reliability of explanations.
@@
$\bullet$ We motivate the need for modeling uncertainty for faithful estimation of concept explanations. $\bullet$ We propose a Bayesian estimation method called U-ACE that is both label-free and models uncertainty in the estimation of concept explanations. $\bullet$ We demonstrate the merits of our proposed method U-ACE through theoretical analysis and empirical evidence on two controlled datasets and two real-world datasets.
@@ 2 Background and Motivation
We denote the model-to-be explained as $f:\mathbb{R}^{D}\rightarrow\mathbb{R}^{L}$ that maps D-dimensional inputs to L labels. Further, we use $f^{[l]}(\mathbf{x})$ to denote $l^{th}$ layer representation space and $f(\mathbf{x})[y]$ for $y\in[1,L]$ as the logit for the label $y$. Given a probe-dataset of examples $\mathcal{D}=\{\mathbf{x}^{(i)}\}_{i=1}^{N}$ and a list of concepts $\mathcal{C}=\{c_{1},c_{2},\dots,c_{K}\}$, our objective is to explain the pretrained model $f$ using the specified concepts. Traditionally, the concepts are demonstrated using potentially small and independent datasets with concept annotations $\{\mathcal{D}_{c}^{k}:k\in[1,K]\}$ where $\mathcal{D}_{c}^{k}$ is a dataset with positive and negative examples of the $k^{th}$ concept.
@@
Concept-Based Explanations (CBE) estimate explanations in two steps. In the first step, they learn what are known as concept activation vectors that predict the concept from $l^{th}$ layer representation of an example. More formally, they learn the concept activation vector $v_{k}$ for $k^{th}$ concept by optimizing $v_{k}=\operatorname*{arg\,min}_{v}\mathbb{E}_{(x,y)\sim\mathcal{D}_{c}^{(k)}}[\ell(v^{T}f^{[l]}(\mathbf{x}),y)]$ where $\ell$ is the usual cross-entropy loss. The inner product of representation with the concept activation vector $v_{k}^{T}f^{[l]}(\mathbf{x})$ is usually referred to as concept activations. Various approaches exist to aggregate example-specific concept activations in to global example-agnostic explanations for the second step. Kim et al. (2018) computes sensitivity of logits to interventions on concept activations to compute what is known as CAV score per example per concept and report the fraction of examples in the probe-dataset with a positive CAV score as the global importance of the concept known as TCAV score. Zhou et al. (2018) proposed to decompose the classification layer weights as $\sum_{k}\alpha_{k}v_{k}$ and report the coefficients $\alpha_{k}$ as the importance score of the $k^{th}$ concept. We refer the reader to Yeh et al. (2022) for an in-depth survey.
@@
Data-efficient concept explanations. A major limitation of traditional CBEs is their need for datasets with concept annotations $\{\mathcal{D}_{c}^{1},\mathcal{D}_{c}^{2},\dots\}$. In practical applications, we may wish to find important concepts among thousands of potentially relevant concepts, which is not possible without expensive data collection. Recent proposals (Yuksekgonul et al., 2022; Oikarinen et al., 2023; Moayeri et al., 2023) suggested using pretrained multi-modal models like CLIP to evade the data annotation cost for a related problem called Concept Bottleneck Models (CBM) (Koh et al., 2020). CBMs aim to train inherently interpretable model with a concept bottleneck. Although CBMs cannot generate explanations for a model-to-be-explained, a subset of methods propose to train what are known as Posthoc-CBMs using the representation layer of a pretrained task model for data efficiency. Given that Posthoc-CBMs base on the representation of a pretrained task model, we may use them to generate concept explanations. We describe briefly two such CBM proposals below.
@@
Oikarinen et al. (2023) (O-CBM) estimates the concept activation vectors by learning to linearly project from the representation space of CLIP where the concept is encoded using its text description to the representation space of the model-to-be-explained $f$. It then learns a linear classification model on concept activations and returns the weight matrix as the concept explanation. Based on the proposal of Yuksekgonul et al. (2022), we can also generate explanations by training a linear model to match the predictions of model-to-be-explained directly using the concept activations of CLIP, which we denote by (Y-CBM).
@@
Unreliable Explanations, a limitation. Apart from data inefficiency, concept explanation methods are known to be unreliable. We observed critical reliability concerns with existing CBEs in the same spirit as the challenges raised in Ramaswamy et al. (2022a). As we demonstrate in Section 4.1, concept explanations for the same model-to-be-explained vary with the choice of the probe-dataset or the concept set bringing into question the reliability of explanations.
@@
Figure 1: Our proposed estimator Uncertainty-Aware Concept Explanations
@@ 3 Uncertainty-Aware Concept Explanations
As summarized in the previous section, CBEs rely on concept activations for generating explanations. It is not hard to see that the activation score of a concept cannot be predicted confidently if the concept is hard/ambiguous or if it is not encoded by the model-to-be-explained. The noise in concept activations if not modeled cascades into the next step leading to poor explanations. Moreover, importance of a concept cannot be confidently estimated if it is missing from the probe-dataset, which must be informed to the user through confidence interval on the concept’s estimated importance score. Motivated by the role of uncertainty for trustworthy explanations, we design our estimator.
Our approach has the following steps. (1) Estimate concept activations along with their error interval, (2) Aggregate concept activations and their confidence intervals in to a global concept explanation. We describe the estimation of concept activations and their error given an instance $\mathbf{x}$ denoted as $\vec{m}(\mathbf{x}),\vec{s}(\mathbf{x})$ respectively in Section 3.1. By definition, the true concept activation for a concept $k$ and instance $\mathbf{x}$ is in the range of $\vec{m}(\mathbf{x})\pm\vec{s}(\mathbf{x})$ with a high probability. We describe the estimation of concept explanations in what follows using $\vec{m}(\mathbf{x}),\vec{s}(\mathbf{x})$, which is independent of how they are computed.
@@
$|\vec{w}^{T}\vec{s}(\mathbf{x})|\leq\delta\quad\forall\mathbf{x}\in\mathcal{D}\implies|\vec{w}^{T}\epsilon|\leq\frac{\sum_{\mathbf{x}\in\mathcal{D}}|\vec{w}^{T}\vec{s}(\mathbf{x})|}{N}\leq\delta\text{ where }\epsilon\triangleq\frac{\sum_{x\in\mathcal{D}}\vec{s}(\mathbf{x})}{N}$
$|\vec{w}^{T}\epsilon|\leq\delta,\text{ for some small }\delta>0\text{ with high probability}\implies\vec{w}^{T}\epsilon\epsilon^{T}\vec{w}\approx\vec{w}^{T}\text{diag}(\epsilon\epsilon^{T})\vec{w}\leq\delta^{2}$
$\implies-\frac{1}{2}(\vec{w}-\mathbf{0})^{T}S^{-1}(\vec{w}-\mathbf{0})\text{ where }S^{-1}=\text{diag}(\epsilon\epsilon^{T})\text{ is high when $\vec{w}$ satisfies the constraint}$
$\implies\mathcal{N}(\vec{w};\mathbf{0},\lambda S)\text{ is high for an appropriate $\lambda>0$}\implies\vec{w}\sim\mathcal{N}(\mathbf{0},\lambda S)$
We observe therefore that the weight vectors drawn from $\mathcal{N}(\mathbf{0},\lambda\text{diag}(\epsilon\epsilon^{T})^{-1})$ satisfy the invariance to input noise constraint with high probability. We now estimate the posterior on the weights after having observed the data with the prior on weights set to $\mathcal{N}(0,\lambda\text{diag}(\epsilon\epsilon^{T})^{-1})$. The posterior over weights has the following closed form(Salakhutdinov, 2011) where $C_{X}=[\vec{m}(\mathbf{x}_{1}),\vec{m}(\mathbf{x}_{2}),\dots,\vec{m}(\mathbf{x}_{N})]$ is a $K\times N$ matrix and $Y=[f(\mathbf{x}_{1})[y],f(\mathbf{x}_{2})[y],\dots,f(\mathbf{x}_{N})[y]]^{T}$ is an $N\times 1$ vector (derivation in Appendix A.1).
$\Pr(\vec{w}\mid C_{X},Y)=\mathcal{N}(\vec{w};\mu,\Sigma)\qquad\text{ where }\mu=\beta\Sigma C_{X}Y,\quad\Sigma^{-1}=\beta C_{X}C_{X}^{T}+\lambda^{-1}\text{diag}(\epsilon\epsilon^{T})$ (1)
@@
Sparsifying weights for interpretability. Because a dense weight matrix can be hard to interpret, we induce sparsity in $W_{c}$ by setting all the values below a threshold to zero. The threshold is picked such that the accuracy on train split does not fall by more than $\kappa$, which is a positive hyperparameter.
The estimator shown in Equation 1 and details on how we estimate the noise in concept activations presented in the next section completes the description of our estimator. We call our estimator Uncertainty-Aware Concept Explanations (U-ACE) because it computes and models the uncertainty in concept activations. Algorithm 1 summarizes our proposed system.
@@ 3.1 Estimation of concept activations and their noise
In this section, we discuss how we estimate $\vec{m}(\mathbf{x}),\vec{s}(\mathbf{x})$ using a pretrained multi-modal model. Recall that image-text multi-modal (MM) systems such as CLIP (Radford et al., 2021) can embed both images and text in a shared representation space, which enables one to estimate the similarity of an image to any phrase. This presents us an interesting solution approach of specifying a concept using its text description ($T_{k}$ for the $k^{th}$ concept) without needing concept datasets $\mathcal{D}_{c}^{k}$. We denote by $g(\bullet)$ the image embedding function of MM and $g_{text}(\bullet)$ the text embedding function.
Our objective is to estimate $\vec{m}(\mathbf{x}),\vec{s}(\mathbf{x})$ such that the true concept activation value is in the range $\vec{m}(\mathbf{x})\pm\vec{s}(\mathbf{x})$. Two major sources of uncertainty in concept activations are due to (1) epistemic uncertainty arising from lack of information about the concept in the representation layer of the model-to-be-explained, (2) data uncertainty arising from ambiguity (because the concept is not clearly visible, see Appendix G.1 for some examples). We wish to estimate $\vec{s}(\mathbf{x})$ that is aware of both the forms of uncertainty.
We can obtain a point estimate for the activation vector of the $k^{th}$ concept $v_{k}$ such that $f(\mathbf{x})^{T}v_{k}\approx g(\mathbf{x})^{T}w_{k}$ (where $w_{k}=g_{text}(T_{k})$) for all $\mathbf{x}$ in the probe-dataset $\mathcal{D}$ through simple optimization (Oikarinen et al., 2023; Moayeri et al., 2023). We may then simply repeat the estimation procedure multiple times to sample from the distribution of activation vectors and their corresponding concept activations. However, as shown empirically in Appendix G.1, $\vec{s}(\mathbf{x})$ estimated from random sampling is a poor measure of uncertainty.
@@
Borrowing from Oikarinen et al. (2023), we define $cos(\alpha_{k})$ as
$\max_{v}[\text{cos-sim}(e(v,f,X),e(w_{k},g,\mathcal{D}))]$ where $e(w_{k},g,\mathcal{D})\triangleq[w_{k}^{T}g(\mathbf{x}_{1}),\dots,w_{k}^{T}g(\mathbf{x}_{N})]^{T}$, and
@@
The proof can be found in Appendix C. The mean and scale values above have a clean interpretation. If the model-to-be-explained ($f$) uses the $k^{th}$ concept for label prediction, the information about the concept is encoded in $f$ and we get a good fit, i.e. $cos(\alpha_{k})\approx 1$, and a small error on concept activations. On the other hand, error bounds are large and concept activations are suppressed when the fit is poor, i.e. $cos(\alpha_{k})\approx 0$. In Appendix G.1, we contrasted different methods for estimation of $\vec{s}(\mathbf{x})$. We observed from the empirical evaluation that U-ACE modeled both model and data uncertainty well.

### 05 Apr 2024 v2

# Estimation of Concept Explanations Should be Uncertainty-Aware

@@ Abstract
Model explanations can be valuable for interpreting and debugging predictive models. We study a specific kind called Concept Explanations, where the goal is to interpret a model using human-understandable concepts. Although popular for their easy interpretation, concept explanations are known to be noisy. We begin our work by identifying various sources of uncertainty in the estimation pipeline that lead to such noise. We then propose an uncertainty-aware Bayesian estimation method to address these issues, which readily improved the quality of explanations. We demonstrate with theoretical analysis and empirical evaluation that explanations computed by our method are robust to train-time choices while also being label-efficient. Further, our method proved capable of recovering relevant concepts amongst a bank of thousands, in an evaluation with real-datasets and off-the-shelf models, demonstrating its scalability. We believe the improved quality of uncertainty-aware concept explanations make them a strong candidate for more reliable model interpretation. We release our code at https://github.com/vps-anonconfs/uace.
@@ 1 Introduction
As increasingly complex machine learning (ML) systems proliferate into real-world decision-making and reasoning, there is a growing need to be able to explain such systems. Concept-based explanations are a class of interpretable methods that explain predictions using high-level and semantically meaningful concepts [Kim et al., 2018]. Concept explanations are aligned with how humans communicate their decisions [Yeh et al., 2022] and are shown [Kim et al., 2018, 2023b] to be more preferable over explanations using salient input features [Ribeiro et al., 2016, Selvaraju et al., 2017] or salient training examples [Koh and Liang, 2017]. Concept explanations also show promise for encoding task-specific prior knowledge [Yuksekgonul et al., 2022] and propelling scientific discovery [Yeh et al., 2022].
@@
Concept explanations offer insight into a pretrained prediction model by estimating the importance of concepts using two human-provided resources: (1) a list of potentially relevant concepts for the task, and (2) a dataset of examples, often referred to as the “probe-dataset”. Estimation typically proceeds in two steps: (a) compute the log-likelihood of a concept given an example called concept activations, and (b) aggregate their per-example activation scores into a globally relevant explanation. For example, the concept wing is considered important if the information about the concept is encoded in all examples of the plane class in the dataset. Owing to their model-level granularity, concept explanations are easy to interpret and have witnessed wide recognition in diverse applications [Yeh et al., 2022].
@@
Notwithstanding their popularity, concept explanations are known to be noisy and data expensive. Ramaswamy et al. [2022a] showed that existing estimation methods are sensitive to the choice of concept set and dataset raising concerns over their interpretability. Another major limitation of concept-based explanation is the need for datasets with explicit concept annotations. Increasingly popular multimodal models such as CLIP [Radford et al., 2021] present an exciting alternate direction to specify relevant concepts, especially for common image applications, through their text description. Recent work has explored using multimodal models for training Concept Bottleneck Models (CBMs) [Oikarinen et al., 2023, Yuksekgonul et al., 2022, Moayeri et al., 2023], but they are not yet evaluated for generating post-hoc concept explanations.
@@
Our objective is to improve the quality of concept explanations while also not requiring datasets with concept annotations. We begin by observing that existing estimation methods do not model uncertainty in the estimation pipeline, leading to high variance or noisy explanations. We identify at least two sources of uncertainty in the standard estimation of concept explanations. First, when a concept is missing from the probe-dataset, we cannot estimate its importance with confidence. Reporting uncertainty over estimated importance of a concept can thus help the user draw a more informed interpretation. Second, when a concept is hard to recognize, or irrelevant to the task, the corresponding activations predicted from the representation layer of the model-to-be-explained are expected to be noisy. If not modelled, uncertainty over concept activations either due to their absence, hardness, or relevance cascades into noise in explanations. Appreciating the need to model uncertainty, we present an estimator called Uncertainty-Aware Concept Explanations (U-ACE), which we show is instrumental in improving quality of explanations.
@@
$\bullet$ We motivate the need for modelling uncertainty for faithful estimation of concept explanations. $\bullet$ We propose a Bayesian estimation method called U-ACE that is both label-free and models uncertainty in the estimation of concept explanations. $\bullet$ We demonstrate the merits of our proposed method U-ACE through theoretical analysis and empirical evidence on two controlled and three real-world datasets.
@@ 2 Background and Motivation
We denote the model-to-be explained as $f:\mathbb{R}^{D}\rightarrow\mathbb{R}^{L}$ that maps D-dimensional inputs to L labels. Let $f^{[l]}(\mathbf{x})$ denote the $l^{th}$ layer representation space and $f(\mathbf{x})[y]$ for $y\in[1,L]$ be the logit for the label $y$. Given a probe-dataset of examples $\mathcal{D}=\{\mathbf{x}^{(i)}\}_{i=1}^{N}$ and a list of concepts $\mathcal{C}=\{c_{1},c_{2},\dots,c_{K}\}$, our objective is to explain the pretrained model $f$ using the specified concepts. Traditionally, the concepts are demonstrated using potentially small and independent datasets with concept annotations $\{\mathcal{D}_{c}^{k}:k\in[1,K]\}$ where $\mathcal{D}_{c}^{k}$ is a dataset with positive and negative examples of the $k^{th}$ concept.
@@
Concept-Based Explanations (CBE) proceed in two steps. In the first step, they learn concept activation vectors (CAVs) that predict the concept from $l^{th}$ layer representation of an example. More formally, they learn the concept activation vector $v_{k}$ for $k^{th}$ concept by optimizing $v_{k}=\operatorname*{arg\,min}_{v}\mathbb{E}_{(x,y)\sim\mathcal{D}_{c}^{(k)}}[\ell(v^{T}f^{[l]}(\mathbf{x}),y)]$ where $\ell$ is the usual cross-entropy loss. The inner product of the representation with the concept activation vector $v_{k}^{T}f^{[l]}(\mathbf{x})$ is usually referred to as the concept activations. Various approaches exist to aggregate example-specific concept activations into model-level explanations for the second step. Kim et al. [2018] computes sensitivity of logits to interventions on concept activations to compute what is known as the CAV score per example per concept and report the fraction of examples in the probe-dataset with a positive CAV score as the global importance of the concept known as TCAV score. Zhou et al. [2018] proposed to decompose the classification layer weights as $\sum_{k}\alpha_{k}v_{k}$ and report the coefficients $\alpha_{k}$ as the importance score of the $k^{th}$ concept. We refer the reader to Yeh et al. [2022] for an in-depth survey.
@@
Data-efficient concept explanations. A major limitation of traditional CBEs is their need for datasets with concept annotations $\{\mathcal{D}_{c}^{1},\mathcal{D}_{c}^{2},\dots\}$. In practical applications, we may wish to find important concepts among thousands of potentially relevant concepts, which is not possible without expensive data collection. Recent proposals [Yuksekgonul et al., 2022, Oikarinen et al., 2023, Moayeri et al., 2023] suggested using pretrained multimodal models like CLIP to evade the data annotation cost for a related problem called Concept Bottleneck Models (CBM) [Koh et al., 2020]. CBMs aim to train inherently interpretable model with a concept bottleneck. Although CBMs cannot generate explanations for a model-to-be-explained, a subset of methods propose to train what are known as Posthoc-CBMs using the representation layer of a pretrained task model for data efficiency. Given that Posthoc-CBMs base on the representation of a pretrained task model, we may use them to generate concept explanations. We describe briefly two such CBM proposals below.
@@
Oikarinen et al. [2023] (O-CBM) estimates the concept activation vectors by learning to linearly project from the representation space of CLIP where the concept is encoded using its text description to the representation space of the model-to-be-explained $f$. It then learns a linear classification model on concept activations and returns the weight matrix as the concept explanation. Based on the proposal of Yuksekgonul et al. [2022], we can also generate explanations by training a linear model to match the predictions of model-to-be-explained directly using the concept activations of CLIP, which we denote by (Y-CBM).
@@
Noisy explanations. Alongside data inefficiency, concept explanation methods are known to be noisy. We observed critical concerns with existing CBEs in the same spirit as the challenges raised in Ramaswamy et al. [2022a]. As we demonstrate in Sections 56.1G.1, concept explanations for the same model-to-be-explained vary with the choice of the probe-dataset or the concept set bringing into question the reliability of explanations.
@@
Figure 1: Our proposed estimator: Uncertainty-Aware Concept Explanations (U-ACE). We track uncertainty in concept activation scores from Step 1, and model them in Step 2.
@@ 3 Uncertainty-Aware Concept Explanations
As summarized in the previous section, CBEs rely on concept activations for generating explanations. It is not hard to see that the activation score of a concept cannot be predicted confidently if the concept is hard/ambiguous, or if it is not encoded by the model-to-be-explained. Moreover, the importance of a concept cannot be confidently estimated if it is missing from the probe-dataset. Motivated by the need to model uncertainty in the estimation, we design our estimator: U-ACE.
Our approach encompasses the following steps: (1) estimate concept activations along with their error interval, and (2) aggregate concept activations and their confidence intervals in to a global concept explanation. We describe the estimation of concept activations and their error given an instance $\mathbf{x}$ denoted as $\vec{m}(\mathbf{x}),\vec{s}(\mathbf{x})$ respectively in Section 3.1. By definition, the true concept activation for a concept $k$ and instance $\mathbf{x}$ is in the range of $\vec{m}(\mathbf{x})\pm\vec{s}(\mathbf{x})$ with a high probability. We describe the estimation of concept explanations in what follows using $\vec{m}(\mathbf{x}),\vec{s}(\mathbf{x})$, which is independent of how they are computed.
@@
The per-example constraint $|\vec{w}^{T}\vec{s}(\mathbf{x})|\leq\delta$ leads to the following constraint with the average noise.
$|\vec{w}^{T}\epsilon|\leq\frac{\sum_{\mathbf{x}\in\mathcal{D}}|\vec{w}^{T}\vec{s}(\mathbf{x})|}{N}\leq\delta\text{ where }\epsilon\triangleq\frac{\sum_{x\in\mathcal{D}}\vec{s}(\mathbf{x})}{N}$
If we require that the dot-product with the average noise be bounded, i.e. $|\vec{w}^{T}\epsilon|\leq\delta,\text{ for some small }\delta>0$ with high probability, their norm, $\vec{w}^{T}\epsilon\epsilon^{T}\vec{w}$, must be bounded. If we then approximate $\epsilon\epsilon^{T}$ with the positive semi-definite $\text{diag}(\epsilon\epsilon^{T})$, we can pose the constraint as a distributional prior as below.
$\vec{w}^{T}\text{diag}(\epsilon\epsilon^{T})\vec{w}\leq\delta^{2}$
$\Rightarrow-\frac{1}{2}(\vec{w}-\mathbf{0})^{T}S^{-1}(\vec{w}-\mathbf{0})\text{ where }S^{-1}=\text{diag}(\epsilon\epsilon^{T})$
is high when $\vec{w}$ satisfies the constraint
$\Rightarrow\mathcal{N}(\vec{w};\mathbf{0},\lambda S)\text{ is high for an appropriate $\lambda>0$}$
$\Rightarrow\vec{w}\sim\mathcal{N}(\mathbf{0},\lambda S)$
We observe therefore that the weight vectors drawn from $\mathcal{N}(\mathbf{0},\lambda\text{diag}(\epsilon\epsilon^{T})^{-1})$ satisfy the invariance to input noise constraint with high probability. We now estimate the posterior on the weights after having observed the data with the prior on weights set to $\mathcal{N}(0,\lambda\text{diag}(\epsilon\epsilon^{T})^{-1})$. The posterior over weights has the following closed form[Salakhutdinov, 2011] where $C_{X}=[\vec{m}(\mathbf{x}_{1}),\vec{m}(\mathbf{x}_{2}),\dots,\vec{m}(\mathbf{x}_{N})]$ is a $K\times N$ matrix and $Y=[f(\mathbf{x}_{1})[y],f(\mathbf{x}_{2})[y],\dots,f(\mathbf{x}_{N})[y]]^{T}$ is an $N\times 1$ vector (derivation in Appendix A.1).
$\Pr(\vec{w}\mid C_{X},Y)=\mathcal{N}(\vec{w};\mu,\Sigma)$ (1)
[…]
@@
The estimator shown in Equation 1 is how we model noise in U-ACE. In the next section, we describe how we may estimate the noise in the concept activations. Algorithm 1 summarizes our proposal.
@@ 3.1 Estimation of concept activations and their noise
In this section, we discuss how we estimate $\vec{m}(\mathbf{x}),\vec{s}(\mathbf{x})$ using a pretrained multimodal model. Recall that image-text multimodal (MM) systems such as CLIP [Radford et al., 2021] can embed both images and text in a shared representation space, which enables one to estimate the similarity of an image to any phrase. This presents us with the flexibility to specify a concept using its text description ($T_{k}$ for the $k^{th}$ concept) without needing concept datasets $\mathcal{D}_{c}^{k}$. We denote by $g(\bullet)$ the image embedding function of MM and $g_{text}(\bullet)$ the text embedding function.
Our objective is to estimate $\vec{m}(\mathbf{x}),\vec{s}(\mathbf{x})$ such that the true concept activation value is in the range $\vec{m}(\mathbf{x})\pm\vec{s}(\mathbf{x})$. Two major sources of uncertainty in concept activations that must inform $s(\mathbf{x})$ are due to (1) epistemic (model) uncertainty arising from lack of information about the concept in the representation layer of the model-to-be-explained, (2) data uncertainty arising from ambiguity (because the concept is not clearly visible, see Figures 7, 8, 9, 10 of Appendix H.1 for some examples). We wish to estimate $\vec{s}(\mathbf{x})$ that is aware of both the forms of uncertainty.
We can obtain a point estimate for the activation vector of the $k^{th}$ concept $v_{k}$ such that $f(\mathbf{x})^{T}v_{k}\approx g(\mathbf{x})^{T}w_{k}$ (where $w_{k}=g_{text}(T_{k})$) for all $\mathbf{x}$ in the probe-dataset $\mathcal{D}$ through simple optimization [Oikarinen et al., 2023, Moayeri et al., 2023]. We may then simply repeat the estimation procedure multiple times to sample from the distribution of activation vectors and their corresponding concept activations. However, as shown empirically in Appendix H.1, $\vec{s}(\mathbf{x})$ estimated from random sampling is a poor measure of uncertainty, which is unsurprising for high dimensional spaces.
@@
Borrowing from Oikarinen et al. [2023], we define $cos(\alpha_{k})$
as $\max_{v}[\text{cos-sim}(e(v,f,\mathcal{D}),e(w_{k},g,\mathcal{D}))]$ where $e(w_{k},g,\mathcal{D})\triangleq[w_{k}^{T}g(\mathbf{x}_{1}),\dots,w_{k}^{T}g(\mathbf{x}_{N})]^{T}$, and
@@
The proof can be found in Appendix C. The mean and scale values above have a clean interpretation. If the model-to-be-explained ($f$) uses the $k^{th}$ concept for label prediction, the information about the concept is encoded in $f$ and we get a good fit, i.e. $cos(\alpha_{k})\approx 1$, and a small error on concept activations. On the other hand, error bounds are large and concept activations are suppressed when the fit is poor, i.e. $cos(\alpha_{k})\approx 0$.
Although $s(\mathbf{x})$ given by Proposition 1 is simple, we found it to be surprisingly effective, which are presented in Appendix H.1. The appendix section contains the following details. (1) We introduced two other variants for measuring uncertainty: Monte Carlo and a distribution fit method (based on [Kim et al., 2023a]). (2) We evaluated different methods for quantifying the different sources of uncertainty on a real dataset. Across different sources of uncertainty and uncertainty estimation methods, we found our method to be consistently effective.